Yellowstone Mining Company had total depletable capitalized costs of $684,000 for a mine acquired in early 2019. It was estimated that the mine contained 760,000 tons of recoverable ore when production began. During 2019, 38,000 tons were mined, and 144,400 tons were mined in 2020.

Required:

  1. 1. Compute the depletion expense in 2019 and 2020 for financial accounting purposes.

  2. 2-a. In 2019, 38,000 tons of ore were sold for $3,800,000. For tax purposes, operating expenses of the mine were $440,000. The taxpayer may deduct either cost depletion or percentage depletion, which for the type ore produced is 8 percent of production sold from the mine. (Assume, however, that percentage depletion is limited to the amount of net income from the property.) What would be the amount of percentage depletion allowable in 2019?

  3. 2-b. What would be the amount of cost depletion allowable for tax purposes in 2019, assuming that capitalized mineral costs are the same for tax purposes as for financial accounting purposes?

  4. 2-c. What will be the amount of depletion based on cost that the company could deduct on its tax return in 2020 if it deducts percentage depletion in 2019?

  5. 2-d. Suppose that in the first three years of the mine’s life, the company took percentage depletion totaling $680,000. In the fifth year of the mine’s life, production proceeds were $3,700,000. How much percentage depletion could the company deduct in the fifth year, considering allowable percentage depletion rate is 8%?

Complete this question by entering your answers in the tabs below. Required 1 Required 2A Required 2B Required 2C Required 2DRequired 1 Required 2A Required 2B Required 2C Required 2D ................... In 2019, 38,000 tons of ore were sold for $3,8Required 1 Required 2A Required 2B Required 2C Required 2D ................... What would be the amount of cost depletion allRequired 1 Required 2A Required 2B Required 2C Required 2D What will be the amount of depletion based on cost that the companRequired 1 Required 2A Required 2B Required 20 Required 2D Suppose that in the first three years of the mines life, the comp
Answer #1

2019:-684000/760000 = 0.9 per ton = 0.9*38000 = 34200 2020:-0.9*144400 = 129960 2a) percentage depletion in 2019 Sales 3,800,2c) # Particulars Capitalized costs of minerals Less Depletion deducted in 2019 Remaining cost eligible for cost depletion De